A train consists of a caboose (mass = 1000 kg), a
Answer #1

Mass of caboose = m3 = 1000 kg

Mass of middle car = m2 = 2000 kg

Mass of engine car = m1 = 2000 kg

Acceleration of engine = a = 5 m / s2

The engine is pulling the cars in right forward direction. Hence, the tension would be in the left direction.

Tension between middle car and engine car

T = (m2 + m3) * a = 3000 * 5 = 15000 N